OpenStudy (kaylak12345):

5+square root of 3 5-square root of 3

OpenStudy (kaylak12345):

@CGGURUMANJUNATH

OpenStudy (kaylak12345):

they have to be multiplied togther

OpenStudy (studygurl14):

Like this? \(\Large 5+\sqrt{3}( 5-\sqrt{3})\)

OpenStudy (studygurl14):

Or like this? \(\Large (5+\sqrt{3})(5-\sqrt{3})\)

OpenStudy (kaylak12345):

b

OpenStudy (kaylak12345):

the 2nd one

OpenStudy (studygurl14):

Okay. So here's a handy rule... \(\Large a(+b)(a-b) = a^2-b^2\) In this case, a=5, and b=\(\sqrt{3}\)

OpenStudy (studygurl14):

Oops, sorry. \(\Large (a+b)(a-b)=a^2-b^2\)

OpenStudy (kaylak12345):

is the answer 22 or with a radical

OpenStudy (kaylak12345):

@Tlynn514

OpenStudy (studygurl14):

22 is correct. Excellent job!

OpenStudy (kaylak12345):

it is????

OpenStudy (studygurl14):

yep